Functionality Checking

07:58
  • Doing OSCO
08:10
  • Having trouble applying NCPA
  • There seems to be an issue with the pupils
  • There is a bite in the edge which, upon applying NCPA, starts to grow until we get safe skips.
08:33
  • We rotated by 90 degrees.
  • We still see the bite at 10 o'clock.
  • We might need to create some new reconstructors.
  • We can possibly make a new flat?
08:38
  • OCAM not live and the adsec ripped and the adsec went into failure.
AdSecArbitrator _75938|ERR| 90|2022-01-21 08:45:53.278666| IDL > Command returns error : Timeout error
AdSecArbitrator_75938|INF| 91|2022-01-21 08:45:53.278874| COMMANDHANDLER > Command execution failed: Fsm forced to Failure
AdSecArbitrator_75938|ERR| 92|2022-01-21 08:45:53.279043| MAIN > SetFlatAo: FATAL - Timeout error
idl_75860 |INF| 289|2022-01-21 08:45:53.278413| MAIN > Reference wave parameters initialization failed!
AdSecArbitrator_75938|ERR| 93|2022-01-21 08:45:58.480319| HOUSEKEEPER > [AOException] Error (Timeout error) in GetDiagnValue [2] (code -5001) Timeout errorAdSecArbitrator_75938|ERR|

90|2022-01-21 08:45:53.278666| IDL

08:54
  • Everything is back
  • re-sending the preset
09:03
  • saved a new flat <today's date>090307
  • Looking at the out of focus reflection on LUCI, it looks like we may have some vignetting.
  • "default.sav" was pointing to a 2020 flat.
  • 20220121_091715
  • NCPA starting 20220121_091909
  • 20220121_092053
    20220121_092105
    20220121_092117
    20220121_092349
    20220121_092401
    20220121_092414
09:26
  • Going to bin2, yea brother.
09:52
  • GOPT went to 0, for some odd reason, though the process was still running.
09:57
  • Trying again.
  • 20220121_101633 No NCPA
  • 20220121_101824 NCPA
10:28
  • A couple of safe skips when we returned from a large offset.
10:41
  • Heading to bin3
10:48
  • Adsec went into failure.
10:57
  • Re-sending the preset.
  • We were in AUTO and the optimize gain failed
11:10
  • Going
  • Optimize gain runs HO1 to silly values and it fails nearly every time.
  • Manual gains 2.3 1.0 0.0
  • Those are good values, according to the graphs.
  • 20220121_111424
  • The image is elongated.
  • Tweaking the gain doesn't seem to help
  • 20220121_112132
11:22
  • Trying another bin-3 target, to see if we are doing something strange with tip/tilt.
  • Optimize gain failed again with safe skip. It sends the HO1 gain to something like 12, when the optimum value is around 1
  • Something is really wrong with bin3. Tip/tilt is either being badly over-driven or under-driven.
  • slope rms doesn't change, even when I do large changes in tip/tilt
11:49
  • We re-sent the preset at a different PA an optimize gain worked.
  • TT gain is 5.31 HO1 is 0.74
  • We still are having a very elongated image.
  • 20220121_115115
  • The elongation seems fixed WRT the rotation angle.
11:55
  • We're going to fake the system by doing a bin2 star but telling it to go to bin3.
  • Now let's see if it's a banana
  • it started round and became a banana
  • Something is wrong with bin 3. we get round images then we get bananas
  • 20220121_121152
12:13
  • Heading to bin4
  • 20220121_123421
  • It's "round-ish"
  • 20220121_123616
  • GOPT is not entirely stable.
  • 20220121_125215 No GOPT
13:05
  • Going to do an ESM preset
  • 20220121_131942
13:35
  • Looks like we're done. We've been guiding on sky brightness.

-- GregoryTaylor - 21 Jan 2022
Topic revision: r1 - 21 Jan 2022, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback